# Decoupled Generative Modeling for Human-Object Interaction Synthesis

### v1

@@ 3.4 Training and Inference
Inference-time Guidance. Inspired by previous work [li2024controllable], our model applies reconstruction guidance to regularize the generation process. This design injects constraints at each denoising step without retraining the network for a specific purpose, which permits flexible control over the outputs. The process is formally represented as:
@@ 4.1 Datasets
Our experiments utilize two datasets designed to capture realistic human-object interactions across diverse indoor environments, following the evaluation setup in prior work [li2024controllable]. i) FullBodyManipulation [li2023object]: Comprising about 10 hours of synchronized human-object motion sequences, this dataset covers 15 distinct rigid objects. We consider only rigid body interactions and therefore exclude sequences with articulated objects. Motion data from 15 subjects are used for training, with data from two additional subjects reserved for testing, ensuring a consistent evaluation protocol. ii) 3D-FUTURE [fu20213d]: Containing a broad range of 3D furniture models. This dataset offers substantial geometric diversity. To evaluate generalization to unseen appearances and shapes, we substitute the 17 test objects in FullBodyManipulation with unseen 3D-FUTURE models of the same categories, enabling systematic assessment on previously unobserved instances.
@@ 4.2 Evaluation Metrics
We evaluate realism and plausibility with five metrics: foot height ($H_{\text{feet}}$), foot sliding ($FS$), R-precision ($R_{\text{prec}}$), Fréchet Inception Distance (FIDFID[heusel2017gans], and Diversity ($DIV$). $H_{\text{feet}}$ is the mean foot-to-ground height. $FS$ measures horizontal foot displacement during stance, following [li2024controllable]. Both are reported in centimeters. $R_{\text{prec}}$ (top-3) scores text-motion alignment. $FID$ compares generated and ground-truth motion distributions in a learned feature space. $DIV$ quantifies variation across samples under the same condition.
@@
Figure 4: Qualitative comparison of DecHOI with CHOIS [li2024controllable] and HOIFHLI [wu2025human] on the FullBodyManipulation [li2023object]. DecHOI produces stable contacts, smooth motion, and accurate object trajectories, while prior methods show drift, penetration, or inconsistent coordination between human and object motions.
@@ 4.3 Quantitative Analysis
We quantitatively evaluate our approach against CHOIS [li2024controllable], HOIFHLI [wu2025human], and two OMOMO [li2023object] variants (Lin- and Pred-OMOMO) implemented following CHOIS on two datasets: FullBodyManipulation [li2023object] and 3D-FUTURE [fu20213d]. Because our method does not require specified intermediate waypoints, we evaluate all baselines without intermediate waypoint inputs to ensure a fair comparison. Note that Lin-OMOMO uses ground truth for all object-related signals, so metrics that are not applicable under this setting are reported as (–).
@@
As shown in Tab. 1, our model surpasses all baselines on most metrics. OMOMO [li2023object] variants condition on object pose at every frame, yet under our extremely limited inputs they underperform in overall human motion and interaction quality. Methods relying on sparse object conditioning [li2024controllable, wu2025human] also degrade on contact-related measures once waypoints are removed, as the reduced conditioning increases task complexity and leads to unsynchronized interactions. In contrast, DecHOI simplifies the problem through decoupled modeling, allowing the trajectory generator to produce accurate paths with low condition matching error and to attain competitive DIV with reduced reliance on waypoints. The resulting interactions exhibit higher realism and stronger alignment with instructions, as reflected by improved FID and $R_{prec}$. Although contact accuracy and penetration scores often trade off against each other, our adversarial training achieves balanced performance across both and produces realistic interactions. Further analyses and ablations are provided in the supplementary material.
@@ 4.4 Qualitative Analysis
In Fig. 4, we present qualitative comparisons among CHOIS [li2024controllable], HOIFHLI [wu2025human], and our DecHOI on the FullBodyManipulation [li2023object]. As illustrated in the four scenes, DecHOI consistently produces more stable and realistic interactions. For the largebox scene, CHOIS and HOIFHLI yield unstable object motion, leading to box hovering and hand-object penetration. In contrast, DecHOI maintains a firmly grounded trajectory and cleaner contacts through stable global path modeling. In the plasticbox scene, baselines exhibit severe penetration between the human body and the object, along with misaligned hand contacts. In contrast, our adversarial training on distal joints mitigates these failures and enhances contact realism. For the tripod scene, both struggle with fine manipulation as the object drifts and rotates inconsistently, whereas our decoupled framework suppresses such unsynchronized artifacts and achieves more coherent motion. Lastly, in the woodchair scene, prior works show noticeable mesh intersection and incomplete contact, while DecHOI produces stable contacts and precise spatial alignment between the hand and object.
Figure 5: Qualitative comparison of DecHOI and CHOIS [li2024controllable] on the 3D-FUTURE [fu20213d], showing generalization to unseen objects.
As shown in Fig. 5, we further compare CHOIS and DecHOI on two unseen objects from the 3D-FUTURE  [fu20213d]: largetable and woodchair. CHOIS often fails to adapt to these novel geometries, leading to human-object intersection and motion instability, whereas our method preserves accurate contacts and smooth motion, demonstrating strong generalization to unseen shapes. More qualitative results and analysis are provided in the supplementary material.
@@ 4.5 Long-term Dynamic Planning
To evaluate long-term and dynamic human-object interactions, we introduce DynaPlan, which supports responsive planning and scene-aware interaction generation under multi-agent conditions. We equip both the agent and the moving counterpart with circular influence radii and detect potential collisions when their regions intersect. When a collision is detected, we re-plan with A*, adaptively choosing either a short detour or waiting to maintain goal-directed and collision-free motion while accounting for the future path of the counterpart. To handle uncertainty in that future path, we predict it with a pre-trained trajectory prediction network [mohamed2020social]. We evaluate performance on about 190 long-sequence indoor scenarios. Further details are provided in the supplementary material.
@@
Figure 6: Visualization of training loss landscapes for DecHOI and CHOIS [li2024controllable], demonstrating reduced optimization complexity.
@@
Table 3: Quantitative results between CHOIS [li2024controllable] and DecHOI for responsive long-term interaction synthesis on the DynaPlan.
Tab. 3 summarizes the quantitative results for long-sequence human-object interaction synthesis on the DynaPlan. Compared to CHOIS [li2024controllable], DecHOI achieves lower trajectory errors $T_{s}$ and $T_{e}$, along with reduced instability, resulting in smoother and more stable motion over extended sequences. For interaction quality, DecHOI shows clear improvements across all penetration metrics, including $P_{\text{hand}}$, $P_{\text{body}}$, $P_{\text{o}\rightarrow\text{s}}$, and $P_{\text{h}\rightarrow\text{s}}$. Here, $P_{\text{hand}}$ and $P_{\text{body}}$ measure the penetration depth between the human mesh and manipulated objects, while $P_{\text{o}\rightarrow\text{s}}$ and $P_{\text{h}\rightarrow\text{s}}$ quantify collisions between the object or human and the surrounding static scene. Lower values on most metrics indicate that DecHOI effectively minimizes interpenetration and maintains collision-free trajectories even in cluttered indoor environments.
@@ 4.6 Ablation Study
Table 4: Ablation results for DecHOI on the FullBodyManipulation [li2023object], evaluating the contribution of each component.
As shown in Tab. 4, we perform an ablation study on the FullBodyManipulation [li2023object] to analyze the effects of adversarial training and text conditioning based on cross-attention. The baseline, which uses only the decoupled generative modeling and concatenates text embeddings along the sequence axis following prior work [li2024controllable, wu2025human], achieves moderate performance but suffers from high penetration errors and limited contact consistency. Introducing adversarial training substantially reduces $P_{\text{hand}}$ and $P_{\text{body}}$, indicating improved realism and more stable interactions, although the linguistic alignment measured by $R_{\text{prec}}$ slightly decreases. Conversely, adding cross-attention layers for text conditioning enhances $R_{\text{prec}}$ by effectively capturing semantic intent, but slightly degrades interaction quality due to weaker regularization. When both modules are combined in DecHOI, the model achieves the best overall performance, simultaneously improving $R_{\text{prec}}$, $C_{F1}$, and penetration metrics. These results indicate that adversarial training enhances the robustness of distal joint interactions during object manipulation, and that text conditioning based on cross-attention enables more explicit transfer of semantic information across modalities, leading to stronger alignment.
@@ 4.7 User Study
Figure 8: Stacked horizontal bars showing user preference distributions for DecHOI compared with CHOIS [li2024controllable] and HOIFHLI [wu2025human] across two evaluation criteria: Text Alignment and Interaction Quality.
We conducted a user study to evaluate the perceptual realism and text alignment of DecHOI compared with CHOIS [li2024controllable] and HOIFHLI [wu2025human]. For each comparison, a total of 60 text-scene pairs, 30 per method, were randomly sampled from our full test set, and 200 participants were recruited via Amazon Mechanical Turk (AMT). Each participant viewed two anonymized clips, each lasting about four seconds, in random order and answered two questions: i) which video better matches the text instruction (Text Alignment), and ii) which video shows better human-object interaction quality (Interaction Quality). Responses were collected using a two-choice scale: Prefer Ours or Prefer Others. As summarized in Fig. 8, DecHOI was consistently preferred on both criteria across both baselines, confirming that our decoupled framework yields motions that are semantically faithful to textual instructions and more realistic, stable, and natural than prior methods.

### v2

@@ 3.4 Training and Inference
Inference-time Guidance. Inspired by previous work [23], our model applies reconstruction guidance to regularize the generation process. This design injects constraints at each denoising step without retraining the network for a specific purpose, which permits flexible control over the outputs. The process is formally represented as:
@@ 4.1 Datasets
Our experiments utilize two datasets designed to capture realistic human-object interactions across diverse indoor environments, following the evaluation setup in prior work [23]. i) FullBodyManipulation [24]: Comprising about 10 hours of synchronized human-object motion sequences, this dataset covers 15 distinct rigid objects. We consider only rigid body interactions and therefore exclude sequences with articulated objects. Motion data from 15 subjects are used for training, with data from two additional subjects reserved for testing, ensuring a consistent evaluation protocol. ii) 3D-FUTURE [9]: Containing a broad range of 3D furniture models. This dataset offers substantial geometric diversity. To evaluate generalization to unseen appearances and shapes, we substitute the 17 test objects in FullBodyManipulation with unseen 3D-FUTURE models of the same categories, enabling systematic assessment on previously unobserved instances.
@@ 4.2 Evaluation Metrics
We evaluate realism and plausibility with five metrics: foot height ($H_{\text{feet}}$), foot sliding ($FS$), R-precision ($R_{\text{prec}}$), Fréchet Inception Distance (FIDFID[14], and Diversity ($DIV$). $H_{\text{feet}}$ is the mean foot-to-ground height. $FS$ measures horizontal foot displacement during stance, following [23]. Both are reported in centimeters. $R_{\text{prec}}$ (top-3) scores text-motion alignment. $FID$ compares generated and ground-truth motion distributions in a learned feature space. $DIV$ quantifies variation across samples under the same condition.
@@
Figure 4: Qualitative comparison of DecHOI with CHOIS [23] and HOIFHLI [43] on the FullBodyManipulation [24]. DecHOI produces stable contacts, smooth motion, and accurate object trajectories, while prior methods show drift, penetration, or inconsistent coordination between human and object motions.
@@ 4.3 Quantitative Analysis
We quantitatively evaluate our approach against CHOIS [23], HOIFHLI [43], and two OMOMO [24] variants (Lin- and Pred-OMOMO) implemented following CHOIS on two datasets: FullBodyManipulation [24] and 3D-FUTURE [9]. Because our method does not require specified intermediate waypoints, we evaluate all baselines without intermediate waypoint inputs to ensure a fair comparison. Note that Lin-OMOMO uses ground truth for all object-related signals, so metrics that are not applicable under this setting are reported as (–).
@@
As shown in Tab. 1, our model surpasses all baselines on most metrics. OMOMO [24] variants condition on object pose at every frame, yet under our extremely limited inputs they underperform in overall human motion and interaction quality. Methods relying on sparse object conditioning [23, 43] also degrade on contact-related measures once waypoints are removed, as the reduced conditioning increases task complexity and leads to unsynchronized interactions. In contrast, DecHOI simplifies the problem through decoupled modeling, allowing the trajectory generator to produce accurate paths with low condition matching error and to attain competitive DIV with reduced reliance on waypoints. The resulting interactions exhibit higher realism and stronger alignment with instructions, as reflected by improved FID and $R_{prec}$. Although contact accuracy and penetration scores often trade off against each other, our adversarial training achieves balanced performance across both and produces realistic interactions. Further analyses and ablations are provided in the supplementary material.
@@ 4.4 Qualitative Analysis
In Fig. 4, we present qualitative comparisons among CHOIS [23], HOIFHLI [43], and our DecHOI on the FullBodyManipulation [24]. As illustrated in the four scenes, DecHOI consistently produces more stable and realistic interactions. For the largebox scene, CHOIS and HOIFHLI yield unstable object motion, leading to box hovering and hand-object penetration. In contrast, DecHOI maintains a firmly grounded trajectory and cleaner contacts through stable global path modeling. In the plasticbox scene, baselines exhibit severe penetration between the human body and the object, along with misaligned hand contacts. In contrast, our adversarial training on distal joints mitigates these failures and enhances contact realism. For the tripod scene, both struggle with fine manipulation as the object drifts and rotates inconsistently, whereas our decoupled framework suppresses such unsynchronized artifacts and achieves more coherent motion. Lastly, in the woodchair scene, prior works show noticeable mesh intersection and incomplete contact, while DecHOI produces stable contacts and precise spatial alignment between the hand and object.
Figure 5: Qualitative comparison of DecHOI and CHOIS [23] on the 3D-FUTURE [9], showing generalization to unseen objects.
As shown in Fig. 5, we further compare CHOIS and DecHOI on two unseen objects from the 3D-FUTURE  [9]: largetable and woodchair. CHOIS often fails to adapt to these novel geometries, leading to human-object intersection and motion instability, whereas our method preserves accurate contacts and smooth motion, demonstrating strong generalization to unseen shapes. More qualitative results and analysis are provided in the supplementary material.
@@ 4.5 Long-term Dynamic Planning
To evaluate long-term and dynamic human-object interactions, we introduce DynaPlan, which supports responsive planning and scene-aware interaction generation under multi-agent conditions. We equip both the agent and the moving counterpart with circular influence radii and detect potential collisions when their regions intersect. When a collision is detected, we re-plan with A*, adaptively choosing either a short detour or waiting to maintain goal-directed and collision-free motion while accounting for the future path of the counterpart. To handle uncertainty in that future path, we predict it with a pre-trained trajectory prediction network [31]. We evaluate performance on about 190 long-sequence indoor scenarios. Further details are provided in the supplementary material.
@@
Figure 6: Visualization of training loss landscapes for DecHOI and CHOIS [23], demonstrating reduced optimization complexity.
@@
Table 3: Quantitative results between CHOIS [23] and DecHOI for responsive long-term interaction synthesis on the DynaPlan.
Tab. 3 summarizes the quantitative results for long-sequence human-object interaction synthesis on the DynaPlan. Compared to CHOIS [23], DecHOI achieves lower trajectory errors $T_{s}$ and $T_{e}$, along with reduced instability, resulting in smoother and more stable motion over extended sequences. For interaction quality, DecHOI shows clear improvements across all penetration metrics, including $P_{\text{hand}}$, $P_{\text{body}}$, $P_{\text{o}\rightarrow\text{s}}$, and $P_{\text{h}\rightarrow\text{s}}$. Here, $P_{\text{hand}}$ and $P_{\text{body}}$ measure the penetration depth between the human mesh and manipulated objects, while $P_{\text{o}\rightarrow\text{s}}$ and $P_{\text{h}\rightarrow\text{s}}$ quantify collisions between the object or human and the surrounding static scene. Lower values on most metrics indicate that DecHOI effectively minimizes interpenetration and maintains collision-free trajectories even in cluttered indoor environments.
@@ 4.6 Ablation Study
Table 4: Ablation results for DecHOI on the FullBodyManipulation [24], evaluating the contribution of each component.
As shown in Tab. 4, we perform an ablation study on the FullBodyManipulation [24] to analyze the effects of adversarial training and text conditioning based on cross-attention. The baseline, which uses only the decoupled generative modeling and concatenates text embeddings along the sequence axis following prior work [23, 43], achieves moderate performance but suffers from high penetration errors and limited contact consistency. Introducing adversarial training substantially reduces $P_{\text{hand}}$ and $P_{\text{body}}$, indicating improved realism and more stable interactions, although the linguistic alignment measured by $R_{\text{prec}}$ slightly decreases. Conversely, adding cross-attention layers for text conditioning enhances $R_{\text{prec}}$ by effectively capturing semantic intent, but slightly degrades interaction quality due to weaker regularization. When both modules are combined in DecHOI, the model achieves the best overall performance, simultaneously improving $R_{\text{prec}}$, $C_{F1}$, and penetration metrics. These results indicate that adversarial training enhances the robustness of distal joint interactions during object manipulation, and that text conditioning based on cross-attention enables more explicit transfer of semantic information across modalities, leading to stronger alignment.
@@ 4.7 User Study
Figure 8: Stacked horizontal bars showing user preference distributions for DecHOI compared with CHOIS [23] and HOIFHLI [43] across two evaluation criteria: Text Alignment and Interaction Quality.
We conducted a user study to evaluate the perceptual realism and text alignment of DecHOI compared with CHOIS [23] and HOIFHLI [43]. For each comparison, a total of 60 text-scene pairs, 30 per method, were randomly sampled from our full test set, and 200 participants were recruited via Amazon Mechanical Turk (AMT). Each participant viewed two anonymized clips, each lasting about four seconds, in random order and answered two questions: i) which video better matches the text instruction (Text Alignment), and ii) which video shows better human-object interaction quality (Interaction Quality). Responses were collected using a two-choice scale: Prefer Ours or Prefer Others. As summarized in Fig. 8, DecHOI was consistently preferred on both criteria across both baselines, confirming that our decoupled framework yields motions that are semantically faithful to textual instructions and more realistic, stable, and natural than prior methods.
@@ 5 Conclusion
Acknowledgement
This work was supported by Culture, Sports and Tourism R&D Program through the Korea Creative Content Agency grant funded by the Ministry of Culture, Sports and Tourism (International Collaborative Research and Global Talent Development for the Development of Copyright Management and Protection Technologies for Generative AI, RS-2024-00345025, 46%), the National Research Foundation of Korea(NRF) grant funded by the Korea government(MSIT)(RS-2025-00521602, 50%), Institute of Information & communications Technology Planning & Evaluation (IITP) & ITRC(Information Technology Research Center) grant funded by the Korea government(MSIT) (No.RS-2019-II190079, Artificial Intelligence Graduate School Program(Korea University), 1%; IITP-2025-RS-2024-00436857, 1%), the National Science Foundation, United States under Grant (NSF) Nos. IIS-2413161 and IIS-2504906, the NVIDIA Academic Grant Program using NVIDIA RTX PRO 6000 Max-Q Workstation Edition.